4.1.31 Question Help Each of two parents has the genotype red /brown, which consists of the pair of alleles that determine hair color, and each parent contibutes one of those alieles to a child Asume th叭ifre chid has at least one red alee, tat color wa domnate and the chids hair color w be red. a List the diflerent possible outcomes Assume that these outcomes are equaly ikely b. What is the probability that a child of these parents will have the brown/brown genolype? C. What is the probability that the child will have red hair color? a. List the possible outoomes O A redI browm and brown/red O B. red /red, redj brown brown/red and brown/brow O C. red /red, red/ brown, and brown/brown O D. redIred and brown/browm and then cick Check
Answer #1

a)

option B is correct : red/red . red/brown .brown/red , red/red

b)

P(brown/brown)=(1/2)*(1/2)=1/4

c)

P(red hair)=1-P(brown/brown)=1-1/4=3/4